Solve for ( x, y)
6x - 2y = 16
y = 2x - 4

Answer by stanbon(75887)

6x - 2y = 16
y = 2x - 4
-------------
Substitute y into the 1st to get:
6x-2(2x-4)=16
Divide thru by 2 to get:
3x-2x+4=8
x=4
Subtitute that into 2nd to get:
y=2*4-4=4
SOLUTION:
x=4,y=4
Cheers,
Stan H.